OpenStudy (anonymous):

Solve the differential equation by separation of variables x(dy/dx)=4y

myininaya (myininaya):

get all your x one on side and y on the opposing side so we have xdy=4ydx dy/y=4dx/x

myininaya (myininaya):

integrate both sides

myininaya (myininaya):

lny=4lnx+C

myininaya (myininaya):

y=e^(lnx+C)

myininaya (myininaya):

oops y=e^(4lnx+C)

OpenStudy (anonymous):

The answer is y=cx^4

OpenStudy (anonymous):

Just can't figure out how they got the x^4

myininaya (myininaya):

y=e^(4lnx)e^C=x^4C

OpenStudy (anonymous):

Thx I see thx a lot :)

myininaya (myininaya):

ok